### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) excel in solving text-based mathematical problems, but they struggle with mathematical diagrams since they are primarily trained on natural scene images. For humans, visual aids generally enhance problem-solving, but MLLMs perform worse as information shifts from textual to visual modality. This decline is mainly due to their shortcomings in aligning images and text. To tackle aforementioned challenges, we propose Math-PUMA, a methodology focused on Progressive Upward Multimodal Alignment. This approach is designed to improve the mathematical reasoning skills of MLLMs through a three-stage training process, with the second stage being the critical alignment stage. We first enhance the language model’s mathematical reasoning capabilities with extensive set of textual mathematical problems. We then construct a multimodal dataset with varying degrees of textual and visual information, creating data pairs by presenting each problem in at least two forms. By leveraging the Kullback-Leibler (KL) divergence of next-token prediction distributions to align visual and textual modalities, consistent problem-solving abilities are ensured. Finally, we utilize multimodal instruction tuning for MLLMs with high-quality multimodal data. Experimental results on multiple mathematical reasoning benchmarks demonstrate that the MLLMs trained with Math-PUMA surpass most open-source MLLMs. Our approach effectively narrows the performance gap for problems presented in different modalities.
@@ Introduction
To address this issue, we propose Math-PUMA, a methodology centered around Progressive Upward Multimodal Alignment (PUMA), aimed at enhancing the mathematical reasoning capabilities of MLLMs. Our approach is structured into three distinct stages, with stage 2 serving as the pivotal alignment phase. (1) Stage 1: We train the LLM using a substantial dataset of text-based math problems to enhance its problem-solving capabilities. This phase capitalizes on the extensive availability of text-based math problem-solving data. (2) Stage 2: Through data augmentation leveraging publicly available sources and automated data generation, we have constructed 692K data pairs. Each pair conveys identical information but differs in its multimodal representation. We utilize the KL-divergence of next-token prediction distributions between text-rich and vision-rich problems. This approach ensures the model maintains consistent performance across different modalities, thereby achieving modal alignment and improving its ability to tackle multimodal mathematical problems. (3) Stage 3: We select 996K high-quality multimodal problem-solving data to fine-tune the model, further enhancing its performance in multimodal mathematical problem-solving tasks.
@@
  • Experimental results on three widely-used benchmarks demonstrate that the MLLMs trained with Math-PUMA outperforms most open-source models. Notably, our approach effectively narrows the performance gap for problems that contain the same information but are presented in different modalities, as evidenced by results on MathVerse.
@@
The advancement of MLLMs has driven significant research into multimodal reasoning. Current efforts are primarily centered on data augmentation to improve models’ performance. Significant efforts have been invested in augmenting text-only mathematical problem-solving data to enhance LLMs’ reasoning capabilities (Saxton et al. 2019; Yu et al. 2023; Liu and Yao 2024). G-LLaVA (Gao et al. 2023a) and Math-LLaVA (Shi et al. 2024) improve multimodal mathematical reasoning by constructing the Geo170K and MathV360K datasets, respectively. These datasets are created by generating additional questions for images sourced from public datasets. However, they only serve to expand the text, without increasing the diversity of images in the dataset. GeoGPT4V (Cai et al. 2024) leverages GPT-4V (OpenAI 2023b) to generate new problems and images based on existing datasets, creating a dataset of 4.9K geometric problems combined with 19K open-source data. Nevertheless, due to GPT-4V’s subpar capability in generating code from image descriptions, the quality of the generated data is comparatively inferior. By comparison, our work not only makes new advancements in data augmentation, including text rephrasing and the generation of high-quality images, but also introduces a novel alignment method used for training.
@@
Figure 3: Overview of the PUMA approach. (left) The three stage training process of Math-PUMA. (right) The details for aligning data pair. The input data pair includes text-rich data at the strong level and vision-rich data at the weak level, simultaneously processed by the MLLM. The strong logits and labels are used to supervise the weak logits.
@@
  • Question Design: The Question Designer randomly selects a type of mathematical problem and generates a specific problem based on the selected type. It also randomly selects the information carrier, deciding whether to present the information in text or in an image. This determines the visual information sent to the Plotter and the textual information sent to the Solver.
  • Plotting: Based on the received visual information, the Plotter uses our predefined basic tools to draw diagrams.
  • Problem Solving: The Solver calculates the answer based on the text-only version of the question provided in the received textual information, as this format contains complete problem information. Since the calculation process is executed through a program, the answer is deterministic and reliable. Considering that MLLMs can obtain stronger reasoning abilities from step-by-step solutions, the Solver generates a detailed explanation for each problem by calling GPT-4o-mini (OpenAI 2024a) and verifying the generated explanations against the standard answer to ensure accuracy.
  • Pair Construction: Combining the diagram output from the Plotter and the text output from the Solver, the Pair Constructor can obtain up to four types of data with the same information but different modalities, including: vision-only, vision-dominant, text-dominant, and text-only. Two are randomly selected to form a data pair, with the one containing more information in the visual modality being classified as vision-rich, and the other as text-rich.
@@
We collected 80K mathematical problem-solving data from online sources. By rephrasing the problems from multiple perspectives (Yu et al. 2023), and applying a series of traditional image processing techniques such as scaling, stretching, and gamma transformation, we ultimately expanded the dataset to 310K data. Additionally, we utilized the VisualWebInstruct dataset (TIGER-Lab 2024) containing 262K data. To automate the construction of data pairs, we employed a straightforward text-to-image rendering process to transition the content from text to visual form. The original data serve as the text-rich component, while the generated data form the vision-rich component. In total, we obtained 572K data pairs.
@@
We employ a three-stage pipeline to train our model Math-PUMA, with specific details shown in Figure 3.
@@
Considering the abundance of unsupervised text-based mathematical training corpora and problem-solving data (Shao et al. 2024), in contrast to the scarcity of high-quality multimodal mathematical problem-solving data, we initially train the language model on a large number of text-based math problems to enhance its mathematical reasoning capabilities. Given that some models (Shao et al. 2024; Yang et al. 2024) have been thoroughly trained and have demonstrated superior performance in solving mathematical problems, we respectively use them to initialize our LLM. Subsequently, we extracted 200k data from these datasets (Yue et al. 2023; Tong et al. 2024; Mitra et al. 2024; LI et al. 2024) to fine-tune the model. Through this phase, the language model’s mathematical reasoning abilities are substantially improved.
@@
We observe that the multimodal mathematical reasoning ability of MLLMs resembles a pyramid: from bottom to top, the performance of MLLMs progressively declines while the information in the text modality gradually shifts to the visual modality. To bridge the performance gap between text-rich and vision-rich mathematical reasoning, we propose PUMA, a Progressive Upward Multimodal Alignment method. PUMA aims to enable MLLMs to tackle vision-rich mathematical problems by aligning their outputs with text-rich data, thereby gradually enhancing the reasoning capability of MLLMs on mathematical problems.
@@
$\begin{split}\mathcal{L}_{\text{RKL}}&=\frac{1}{TV}\sum^{T}_{t=1}{\text{KL}\Big{(}p_{i+1}(y^{(i+1)}_{t}|\mathbf{y}^{(i+1)}_{<t})||p_{i}(y^{(i)}_{t}|\mathbf{y}^{(i)}_{<t})\Big{)}}\\ &=\frac{1}{TV}\sum^{T}_{t=1}\sum^{V}_{j=1}p_{i}(Y^{(i+1)}_{j}|\mathbf{y}^{(i+1)}_{<t})\log\frac{p_{i+1}(Y^{(i+1)}_{j}|\mathbf{y}^{(i+1)}_{<t})}{p_{i}(Y^{(i)}_{j}|\mathbf{y}^{(i)}_{<t})},\end{split}$
@@
In the final phase, we enhance the model’s reasoning capabilities by incorporating multimodal problem-solving data. Initially, we retained the majority of the high-quality data used in Stage 2, while augmenting our dataset with the MathV360K dataset (Shi et al. 2024). Specifically, we focused on enriching the geometric problem subset within the MathV360K dataset, expanding it from 40K to 120K data to address the scarcity of geometric data. Additionally, we integrated a balanced amount of textual data to prevent any modal imbalance between text and visual modalities. All data included detailed reasoning processes to guide the model’s understanding and learning.
Ultimately, we compiled a large-scale instruction tuning dataset, comprising a total of 996K data. This multimodal instruction tuning not only bolsters the model’s ability to reason and solve problems but also ensures that it can effectively leverage both textual and visual information for enhanced performance in mathematical problem-solving.
@@
We validate the effectiveness of our method across various base models and scales, including DeepSeek-Math-7B (Shao et al. 2024), Qwen2-1.5B and Qwen2-7B(Yang et al. 2024), chosen as the LLMs for Math-PUMA. To ensure the compatibility with DeepSeek-Math and DeepSeek-VL (Lu et al. 2024a), we adhere to the architecture of DeepSeek-VL. For Qwen2, we adopt a similar architecture to LLaVA, with the visual encoder designated as SigLIP-so400m-patch14-384 (Zhai et al. 2023).
@@
For MathVerse and MathVista, we adopt the official implementation. Initially, we use GPT-4o-mini (OpenAI 2024a) to extract answers from the responses generated by MLLMs. Subsequently, we employ GPT-4o-mini (OpenAI 2024a) once more to verify the correctness of the extracted answers. The prompts used for answer extraction and correctness assessment are kept consistent with the official implementation. Ultimately, we calculate the accuracy scores as the evaluation metric. For We-Math, we select the average and Rote Memorization (RM) scores as evaluation metrics.
@@
Our experiments are conducted using PyTorch version 2.1.0 and CUDA 12.1, utilizing 32 NVIDIA A100 GPUs with 80GB memory each. The training process is divided into three stages, each with specific hyperparameters and configurations. We employ the AdamW optimizer (Kingma and Ba 2014) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is adjusted across three stages: $3\times 10^{-5}$ for stage 1, $5\times 10^{-5}$ for stage 2, and $3\times 10^{-5}$ for stage 3. A cosine learning rate schedule is implemented with a warm-up phase covering 2% of the total training steps. Additionally, a decay rate of 0.1 is applied. The KL divergence is controlled using specific hyperparameters: $\alpha_{\text{KL}}$ is set to 0.2, $\tau$ to 1.0, and $\lambda_{\text{KL}}$ to 0.1. The training is conducted over 1 epoch. The batch sizes for three stages are 256, 512, and 256, respectively.
@@
Close-source LLMs
@@
Close-source MLLMs
@@
MathVerse is capable of clearly demonstrating the gap between visual and textual modalities. From Table 1, it can be observed that the MLLMs trained by Math-PUMA achieve the state-of-the-art (SOTA) among open-source MLLMs. Compared to the previous SOTA method, Math-LLaVA, the MLLMs trained by Math-PUMA exhibit accuracy scores improvement about 10%. When compared to the closed-source GPT-4V (OpenAI 2023b), Math-PUMA-Qwen2-7B performs competitively with only a gap of 6.5%, demonstrating the effectiveness of Math-PUMA.
@@
MathVista is a comprehensive benchmark designed to evaluate mathematical reasoning. According to the results presented in Table 1, Math-PUMA-Qwen2-7B demonstrates SOTA performance in GPS, ALG, GEO and SCI domains among open-source MLLMs of the same scale. It outperforms InternLM-XComposer2-VL (Dong et al. 2024) by significant margins, with accuracy improvements of 16.4%, 15.7%, 16.8%, and 10.7% in these respective domains.
@@
We-Math places strong emphasis on the importance of the mathematical reasoning process. Table 2 demonstrates that Math-PUMA-Qwen2-7B achieves SOTA performance in average scores among open-source MLLMs with approximate 10B parameters, surpassing InternLM-XComposer2-VL (Dong et al. 2024). Notably, even among open-source MLLMs with parameters exceeding 20B, Math-PUMA-Qwen2-7B outperforms LLaVA-NeXT (Liu et al. 2024b) 72B model, reaching the performance of LLaVA-NeXT 110B model. While Math-PUMA surpasses Qwen-VL-Max (Bai et al. 2023) among closed-source models, there remains a significant gap compared to GPT-4V (OpenAI 2023b) and GPT-4o (OpenAI 2024b).
@@
We conduct ablation studies on Mathverse to showcase the effectiveness of each stage and to explore the influence of sequential orders on Math-PUMA.
@@
We design three ablation experiments to demonstrate the effectiveness of each stage by removing stage 1, 2, and 3 separately. Subsequently, we evaluate the accuracy scores of overall, text-dominant and vision-only scenarios, as well as the gap between them. The results of the ablation experiments are presented in Table 3.

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) excel in solving text-based mathematical problems, but they struggle with mathematical diagrams since they are primarily trained on natural scene images. For humans, visual aids generally enhance problem-solving, but MLLMs perform worse as information shifts from textual to visual modality. This decline is mainly due to their shortcomings in aligning images and text. To tackle aforementioned challenges, we propose Math-PUMA, a methodology focused on Progressive Upward Multimodal Alignment. This approach is designed to improve the mathematical reasoning skills of MLLMs through a three-stage training process, with the second stage being the critical alignment stage. We first enhance the language model’s mathematical reasoning capabilities with extensive set of textual mathematical problems. We then construct a multimodal dataset with varying degrees of textual and visual information, creating data pairs by presenting each problem in at least two forms. By leveraging the Kullback-Leibler (KL) divergence of next-token prediction distributions to align visual and textual modalities, consistent problem-solving abilities are ensured. Finally, we utilize multimodal instruction tuning for MLLMs with high-quality multimodal data. Experimental results on multiple mathematical reasoning benchmarks demonstrate that the MLLMs trained with Math-PUMA surpass most open-source MLLMs. Our approach effectively narrows the performance gap for problems presented in different modalities. The code and data are available at: https://github.com/wwzhuang01/Math-PUMA.
@@ Introduction
To address this issue, we propose Math-PUMA, a methodology centered around Progressive Upward Multimodal Alignment (PUMA), aimed at enhancing the mathematical reasoning capabilities of MLLMs. Our approach is structured into three distinct stages, with stage 2 serving as the pivotal alignment phase. (1) Stage 1: We train the LLM using a substantial dataset of text-based math problems to enhance its problem-solving capabilities. This phase capitalizes on the extensive availability of text-based math problem-solving data. (2) Stage 2: It is observed that the model’s mathematical problem-solving ability diminished progressively from text to vision, exhibiting an upward pyramidal structure. Consequently, the model’s capabilities are categorized into four hierarchical levels. We construct 692K data pairs, with each pair conveying identical information but differing in multimodal representation. By leveraging the KL-divergence between next-token prediction distributions for text-rich and vision-rich problems, we achieve progressive bottom-up modal alignment across these hierarchical levels, thereby enhancing the model’s ability to tackle multimodal mathematical problems. (3) Stage 3: We select 996K high-quality multimodal problem-solving data to fine-tune the model, further enhancing its performance in multimodal mathematical problem-solving tasks.
@@
  • Experimental results on three widely-used benchmarks demonstrate that the MLLMs trained with Math-PUMA outperform most open-source models. Notably, our approach effectively narrows the performance gap for problems that contain the same information but are presented in different modalities, as evidenced by results on MathVerse.
@@
The advancement of MLLMs has driven significant research into multimodal reasoning. Current efforts are primarily centered on data augmentation to improve models’ performance. Significant efforts have been invested in augmenting text-only mathematical problem-solving data to enhance LLMs’ reasoning capabilities (Saxton et al. 2019; Yu et al. 2023; Liu and Yao 2024). G-LLaVA (Gao et al. 2023a) and Math-LLaVA (Shi et al. 2024) improve multimodal mathematical reasoning by constructing the Geo170K and MathV360K datasets, respectively. These are created by generating additional questions for images sourced from public datasets. However, they only serve to expand the text, without increasing the diversity of images in the dataset. GeoGPT4V (Cai et al. 2024) leverages GPT-4V (OpenAI 2023b) to generate new problems and images based on existing datasets, creating a dataset of 4.9K geometric problems combined with 19K open-source data. Nevertheless, due to GPT-4V’s subpar capability in generating code from image descriptions, the quality of the generated data is comparatively inferior. By comparison, our work not only makes new advancements in data augmentation, including text rephrasing and the generation of high-quality images, but also introduces a novel alignment method used for training.
@@
Figure 3: Overview of the Math-PUMA approach. (left) The three stage training process of Math-PUMA. (right) The details for aligning data pair. The input data pair includes text-rich data at the strong level and vision-rich data at the weak level, simultaneously processed by the MLLM. The strong logits and labels are used to supervise the weak logits.
@@
  • Question Design: The Question Designer employs a random selection process to determine the specific type of mathematical problem to be generated. It also randomly selects the information carrier, deciding whether to present the information as text or an image. This choice dictates the visual information sent to the Plotter and the textual information sent to the Solver.
  • Plotting: In accordance with the visual information received, the Plotter uses the predefined fundamental tools to plot diagrams.
  • Problem Solving: The Solver calculates the answer using the text-only version of the problem, which contains complete information. As the calculation is performed programmatically, the answer is reliable. Considering that MLLMs can obtain stronger reasoning abilities from step-by-step solutions, the Solver generates a detailed explanation for each problem by calling GPT-4o-mini (OpenAI 2024a) and verifying the explanations against the standard answer to ensure accuracy.
  • Pair Construction: The Pair Constructor combines the diagram from the Plotter and the text from the Solver to obtain up to four types of data, each comprising the same information but presented in a different modality: vision-only, vision-dominant, text-dominant, and text-only. Two of these are randomly selected to form a data pair, with the component containing more visual information classified as vision-rich and the other as text-rich.
@@
We initially collect 80K mathematical problem-solving data from online sources. By rephrasing the problems from multiple perspectives (Yu et al. 2023) and applying a series of traditional image processing techniques such as scaling, stretching, and gamma transformation, we expand the dataset to 310K. Additionally, we utilize the VisualWebInstruct dataset (TIGER-Lab 2024) containing 262K data. To automate the construction of data pairs, we employ a straightforward text-to-image rendering process to convert the content from textual to visual form. The original data serve as the text-rich component, while the generated data form the vision-rich component. In total, we obtain 572K data pairs.
@@
We employ a three-stage pipeline to train our models, with specific details shown in Figure 3.
@@
Given the abundance of unsupervised text-based mathematical training corpora and problem-solving data (Shao et al. 2024), in comparison to the scarcity of high-quality multimodal mathematical problem-solving data, we initially train the LLM on a large corpus of text-based math problems to bolster its mathematical reasoning capabilities. To leverage the strengths of existing LLMs that have demonstrated superior performance in mathematical problem-solving (Shao et al. 2024; Yang et al. 2024), we use them to initialize our MLLMs. Subsequently, we fine-tune the model using 200,000 data points extracted from various datasets (Yue et al. 2023; Tong et al. 2024; Mitra et al. 2024; LI et al. 2024). This phase significantly enhances the LLM’s mathematical reasoning abilities.
@@
We observe that the multimodal mathematical reasoning ability of MLLMs resembles a pyramid, with performance declining from bottom to top as the information shifts from text to visual modalities. In order to address the discrepancy in performance between text-rich and vision-rich mathematical reasoning, we propose PUMA (Progressive Upward Multimodal Alignment). The objective of PUMA is to facilitate the effective resolution of vision-rich mathematical problems by aligning the outputs of MLLMs with text-rich data, thereby enhancing their reasoning capabilities across different modalities.
@@
$\begin{split}\mathcal{L}_{\text{RKL}}&=\frac{1}{TV}\sum^{T}_{t=1}{\text{KL}\Big{(}p_{i+1}(y^{(i+1)}_{t}|\mathbf{y}^{(i+1)}_{<t})||p_{i}(y^{(i)}_{t}|\mathbf{y}^{(i)}_{<t})\Big{)}}\\ &=\frac{1}{TV}\sum^{T}_{t=1}\sum^{V}_{j=1}p_{i+1}(Y^{(i+1)}_{j}|\mathbf{y}^{(i+1)}_{<t})\log\frac{p_{i+1}(Y^{(i+1)}_{j}|\mathbf{y}^{(i+1)}_{<t})}{p_{i}(Y^{(i)}_{j}|\mathbf{y}^{(i)}_{<t})},\end{split}$
@@
In the final phase, we enhance the model’s reasoning capabilities by incorporating multimodal problem-solving data. Initially, we retain the majority of the high-quality data used in Stage 2 and augment our dataset with the MathV360K dataset (Shi et al. 2024). Specifically, we focus on enriching the geometric problem subset within MathV360K, expanding it from 40K to 120K in order to address the scarcity of geometric data. Furthermore, we integrated a balanced amount of textual data to prevent any modal imbalance between text and visual modalities. All data included detailed reasoning processes to guide the model’s understanding and learning.
Ultimately, we compiled a large-scale instruction tuning dataset, comprising a total of 996K data points. This multimodal instruction tuning not only bolsters the model’s reasoning and problem-solving abilities but also ensures that it can effectively leverage both textual and visual information for improved performance in mathematical problem-solving.
@@
We validate the effectiveness of our method across various base models and scales, including DeepSeek-Math-7B (Shao et al. 2024), Qwen2-1.5B and Qwen2-7B(Yang et al. 2024), chosen as the LLM for Math-PUMA. To ensure the compatibility with DeepSeek-Math and DeepSeek-VL (Lu et al. 2024a), we adhere to the architecture of DeepSeek-VL. For Qwen2, we adopt a similar architecture to LLaVA, with the visual encoder designated as SigLIP-so400m-patch14-384 (Zhai et al. 2023).
@@
We refer to the leaderboards and adopt the official implementations of MathVerse, MathVista, and We-Math. For MathVerse and MathVista, initially, we use GPT-4o-mini (OpenAI 2024a) to extract answers from the responses generated by MLLMs. Subsequently, we employ GPT-4o-mini once more to verify the correctness of the extracted answers. The prompts used for answer extraction and correctness assessment are kept consistent with the official implementation. Ultimately, we calculate the accuracy scores as the evaluation metric. For We-Math, we select the average and Rote Memorization (RM) scores as evaluation metrics.
@@
Our experiments are conducted using PyTorch version 2.1.0 and CUDA 12.1, utilizing 32 NVIDIA A100 GPUs with 80GB memory each. The training process is divided into three stages, each with specific hyperparameters and configurations. We employ the AdamW optimizer (Kingma and Ba 2014), configured with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate is adjusted across three stages: $3\times 10^{-5}$ for stage 1, $5\times 10^{-5}$ for stage 2, and $3\times 10^{-5}$ for stage 3. A cosine learning rate schedule is implemented with a warm-up phase covering 2% of the total training steps. Additionally, a decay rate of 0.1 is applied. The KL divergence is controlled using specific hyperparameters: $\alpha_{\text{KL}}$ is set to 0.2, $\tau$ to 1.0, and $\lambda_{\text{KL}}$ to 0.1. The training is conducted over 1 epoch. The batch sizes for three stages are 256, 512, and 256, respectively.
@@
Closed-source LLMs
@@
Closed-source MLLMs
@@
MathVerse is capable of clearly demonstrating the gap between visual and textual modalities. From Table 1, it can be observed that the MLLMs trained by Math-PUMA achieve the state-of-the-art (SOTA) among open-source MLLMs. Compared to the previous SOTA method, Math-LLaVA, the MLLMs trained by Math-PUMA exhibit accuracy scores improvement about 10%. When compared to the closed-source GPT-4V (OpenAI 2023b), Math-PUMA-Qwen2-7B performs competitively with only a gap of 4.7%, demonstrating the effectiveness of Math-PUMA.
@@
MathVista is a comprehensive benchmark designed to evaluate mathematical reasoning. According to the results presented in Table 1, Math-PUMA-Qwen2-7B demonstrates SOTA performance in GPS, ALG, GEO and SCI domains among open-source MLLMs of the same scale. It outperforms InternLM-XComposer2-VL (Dong et al. 2024) by significant margins, with accuracy improvements of 16.4%, 15.7%, 16.8%, and 4.9% in these respective domains.
@@
We-Math places strong emphasis on the importance of the mathematical reasoning process. Table 2 demonstrates that Math-PUMA-Qwen2-7B achieves SOTA performance in average scores among open-source MLLMs with approximate 10B parameters, surpassing InternLM-XComposer2-VL. Notably, even among open-source MLLMs with parameters exceeding 20B, Math-PUMA-Qwen2-7B outperforms LLaVA-NeXT (Liu et al. 2024b) 72B model, reaching the performance of LLaVA-NeXT 110B model. While Math-PUMA-Qwen2-7B surpasses Qwen-VL-Max (Bai et al. 2023) among closed-source models, there remains a significant gap compared to GPT-4V and GPT-4o.
@@
Ablation studies are performed on Mathverse to highlight the contribution of each training stage and to assess the impact of their sequential order on Math-PUMA.
@@
To evaluate the significance of each stage, we conduct three ablation experiments by individually removing stages 1, 2, and 3. We then assess the accuracy across overall, text-dominant, and vision-only scenarios, as well as the gaps between them. The results of these experiments are summarized in Table 3.